Starting from rest,a sprinter reaches his top velocity in 3 seconds.He runs a distance of 24m in the 3 secinds.What is his acceleration?
1
Expert's answer
2019-04-17T11:12:08-0400

We can use the formula for distance assuming constant acceleration:


"s=0.5at^2"

Thus,


"a=\frac{2s}{t^2}=\frac{2(24)}{3^2}=5.3 \frac{m}{s^2}"